Jane started north 2 hours ahead of Julie who was also northbound from the same location. If Julie traveled 60 mph and Jane at 40 mph, how lond did it take Julie to get 20 miles ahead of Jane?
Answer by Paul(988)

Julie started at x hours
Jane started at x+2 hours
40(x+2)-60(x)=20
40x-60x=20-80
-20x=-60
x=3

Hence, it took Julie 3 hours to get 20 milies ahead of Jane.
Paul.